unrise, Inc., has no debt outstanding and a total market value of $240,000. Earnings before interest and taxes, EBIT, are projected to be $28,000 if economic conditions are normal. If there is strong expansion in the economy, then EBIT will be 12 percent higher. If there is a recession, then EBIT will be 25 percent lower. The company is considering a $140,000 debt issue with an interest rate of 6 percent. The proceeds will be used to repurchase shares of stock. There are currently 12,000 shares outstanding. Ignore taxes for this problem. Assume the stock price is constant under all scenarios.

a-1.

Calculate earnings per share (EPS) under each of the three economic scenarios before any debt is issued. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

a-2. Calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
b-1. Calculate earnings per share (EPS) under each of the three economic scenarios assuming the company goes through with recapitalization. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
b-2. Given the recapitalization, calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
a-1
EPS = EBIT*(1-tax rate)/shares outstanding
Recession
EPS = EBIT*(1-recession impact%)*(1-tax rate)/shares outstanding
EPS=28000*(1-0.25)*(1-0)/12000
EPS=1.75
Normal
EPS = EBIT*(1-tax rate)/shares outstanding
EPS=28000*(1-0)/12000
EPS=2.33
Expansion
EPS = EBIT*(1+Growth impact%)*(1-tax rate)/shares outstanding
EPS=28000*(1+0.12)*(1-0)/12000
EPS=2.61
a-2
%age change in EPS for Recession
=(EPS recession/EPS normal-1)*100
=(1.75/2.3333-1)*100
=-25%
%age change in EPS for Growth
=(EPS Growth/EPS normal-1)*100
=(2.6133/2.3333-1)*100
=12%
b-1
New no. of shares = old shares-debt/(Market value/old shares)
=12000-140000/(240000/12000)
=5000
EPS = (EBIT-debt*interest%)*(1-tax rate)/new shares outstanding
Recession
EPS = (EBIT*(1-recession impact%)-debt*interest %age)*(1-tax rate)/new shares outstanding
EPS=(28000*(1-0.25)-140000*0.06)*(1-0)/5000
EPS=2.52
Normal
EPS = (EBIT-debt*interest%)*(1-tax rate)/new shares outstanding
EPS=(28000-140000*0.06)*(1-0)/5000
EPS=3.92
Expansion
EPS = (EBIT*(1+growth impact%)-debt*interest %age)*(1-tax rate)/new shares outstanding
EPS=(28000*(1+0.12)-140000*0.06)*(1-0)/5000
EPS=4.59
b-2
%age change in EPS for Recession
=(EPS recession/EPS normal-1)*100
=(2.52/3.92-1)*100
=-36%
%age change in EPS for Growth
=(EPS Growth/EPS normal-1)*100
=(4.592/3.92-1)*100
=17%